1. A product’s production function is given as Q=50KL. Assume all factors of production are variable. What would be the minimum level of output of production if a firm wants to produce 2000 units of output. Price of labor is 20, price of capital is 80.
1
Expert's answer
2021-12-15T11:35:21-0500

Solution:

Cost minimization:

MRTS = "\frac{MP_{L} }{MP_{K}} = \frac{w }{r}"


MPL = "\frac{\partial Q} {\partial L}" = 50K


MPK = "\frac{\partial Q} {\partial K}" = 50L


"\frac{MP_{L} }{MP_{K}} = \frac{w }{r}"


w = 20

r = 80

"\frac{50K }{50L} = \frac{20 }{80}"


"\frac{K }{L} = 0.25"

K = 0.25L

Q = 50KL

2000 = 50(0.25L) (L) = 12.5L2

2000 = 12.5L2

L2 = 160

L = 12.65

K = 0.25L = 0.25(12.65) = 3.16

The minimum level of output of production (K,L) = (3.16, 12.65)

Labor = 12.65

Capital = 3.16